# X-ray Polarization Detection of the Pulsar Wind Nebula in G21.50.9 with IXPE

### v1

@@ 1 Introduction
G21.50.9 is a young, composite plerionic supernova remnant (SNR) characterized in X-rays by the presence of a central, bright, 40″ radius PWN surrounded by a much dimmer and softer SNR shell (Slane et al., 2000; Safi-Harb et al., 2001). Deep Chandra and XMM-Newton observations have revealed that the SNR blast wave extends up to 150″ and displays a limb brightening feature at the eastern boundary and knots of enhanced soft X-ray emission above the PWN in the northern direction (Matheson & Safi-Harb, 2005; Bocchino et al., 2005; Matheson & Safi-Harb, 2010; Guest et al., 2019). Observations of G21.50.9 across various electromagnetic wavelengths, from radio to -rays, have unveiled a wealth of information about this source. Using the Very Large Array (VLA) to measure its expansion speed, Bietenholz & Bartel (2008) estimated its age to be yr, making G21.50.9 one of the youngest PWNe known in our galaxy. Its distance to Earth was measured by several authors using different techniques and was found to be around 5 kpc (Safi-Harb et al., 2001; Camilo et al., 2006; Bietenholz & Bartel, 2008; Tian & Leahy, 2008). The first maps at radio wavelengths go back to the 1970–80s (Wilson & Weiler, 1976; Becker & Kundu, 1976; Furst et al., 1988), but even more recent deep searches have failed to detect a radio counterpart of the X-ray SNR shell, with the exception of the northern knot (Bietenholz et al., 2011).
At the heart of G21.50.9 lies the pulsar PSR J18331034, first independently discovered in radio by Gupta et al. (2005) and Camilo et al. (2006), who found a compact pulsating source at the center of the PWN (R.A. = , Dec. = (J2000.0)) with period ms, , surface magnetic field of G, characteristic age of 4.8 kyr, and a spin-down luminosity of . The pulsar remains currently undetected at X-ray energies, while -ray pulsation has been detected with the Fermi Large Area Telescope (Abdo et al., 2013; Smith et al., 2023). Past X-ray observations have only led to upper limits on the pulsating fraction, while still detecting a central compact source surrounded by an elliptical emission region (Slane et al., 2000; Safi-Harb et al., 2001; Camilo et al., 2006). The projected spin-axis direction of PSR J18331034, inferred by fitting the pulsar wind torus using Chandra data, was determined to be at a position angle of north through east (Ng & Romani, 2008).
@@
The initial radio polarization maps of G21.50.9 were obtained by Becker & Szymkowiak (1981) with the VLA and by Furst et al. (1988) using the NRO Millimeter-wave Array and revealed a highly ordered, linearly polarized emission distributed in a ring structure, indicating a radial magnetic field configuration. The level of polarization was found to have a minimum towards the center, where little polarization was detected, and to increase with radius up to 20–30% near the boundary of the PWN. Infrared polarimetric observations performed with the Very Large Telescope on the compact core of the PWN revealed a highly polarized emission, with an average value of the linear polarization degree of , and a swing of the polarization angle across the inner nebula consistent with a toroidal magnetic field (Zajczyk et al., 2012). On the other hand, a recent polarimetric study by Lai et al. (2022), using archival high-resolution radio observations taken with VLA, confirmed the global radial magnetic field structure of the PWN and ruled out the existence of any strong large-scale toroidal field components extending beyond the inner nebula.
@@ 2 Observations
IXPE observed G21.50.9 in three separate segments close in time and combined in a single observation with a total livetime of ks: the first segment lasted from October 5, 2023 at 00:05:39.233 UTC to October 9, 2023 at 02:16:33.292 UTC, the second one from October 10, 2023 at 22:45:54.503 UTC to October 17, 2023 at 16:42:55.034 UTC and the last one from October 19, 2023 at 21:35:44.584 UTC to October 27, 2023 at 15:39:00.534 UTC. IXPE Level-2 data, publicly available for download on the HEASARC archive111https://heasarc.gsfc.nasa.gov/docs/ixpe/archive/ (ObsID 02001199), were reprocessed and analyzed according to the standard technique using the version 20240125 of the IXPE Instrument Response Functions (IRFs) released on the CALDB222https://heasarc.gsfc.nasa.gov/docs/ixpe/caldb/ on February 28, 2024 (validity date 20230702). Before analyzing the data, we performed an initial background rejection following the procedure described in Di Marco et al. (2023) to reduce the instrumental background due to cosmic-ray particles interacting in the GPD by approximately 40%. We also removed the short time intervals where sudden spikes in the counting rate are observed connected to solar activity and the South Atlantic Anomaly boundary, and corrected the absolute sky coordinates by adjusting the WCS keywords in the FITS header to align the reference frame to Chandra (TCRPX7 = 304, TCRPX8 = 301.5). These are the data files used from now on in this paper. Fig. 1 shows the count map resulting from the combination of the three IXPE telescopes over the full energy range (2–8 keV). The white circular region (radius 0.8′ centered on the pulsar location) is used to select the inner PWN, while the yellow annulus (inner/outer radii of 3′ and 4′ from the pulsar) defines the background extraction region used for the space-integrated analysis described in Sect. 3.1.
G21.50.9 is observed with Chandra routinely as a calibration source. Overlaid in green in Fig. 1 are the contours of a deep image created by merging 18 individual observations listed in Table 1 (total exposure ks), all of which had the remnant placed in the ACIS-S detector, at the default aimpoint of the S3 CCD chip. The events were combined with the merge_obs script in the ciao333https://cxc.cfa.harvard.edu/ciao/ analysis software. A count map was then created using events in the 0.5–8.0 keV energy band, and adaptively smoothed with the dmimgadapt tool using a cone filter with minimum/maximum smoothing radii of 0.5″ and 150″ spanning 300 logarithmically spaced scales, with a minimum of 75 counts within the smoothing kernel. The resulting map is used as background image in all polarization maps shown in the paper (e.g. Fig. 3).
@@
159 | 14.85 | 3700 | 9.54 |
1230 | 14.56 | 5166 | 10.02 |
1433 | 14.97 | 5159 | 9.83 |
1717 | 7.54 | 6071 | 9.64 |
1770 | 7.22 | 6741 | 9.83 |
1838 | 7.85 | 8372 | 10.01 |
1554 | 9.06 | 10646 | 9.54 |
2873 | 9.83 | 14263 | 9.57 |
3693 | 9.78 | 16420 | 9.57 |
@@ 3.1 Space-integrated analysis
As is typically done for other IXPE observations (Bucciantini et al., 2023b; Xie et al., 2022; Negro et al., 2023), using the spatial regions described at the end of Sect. 2 and shown in Fig. 1, we studied the polarization of the entire PWN in two different ways: a model-independent polarimetric analysis using ixpeobssim (Baldini et al., 2022) and a spectro-polarimetric fit with the Multi-Mission Maximum Likelihood (3ML) framework (Vianello et al., 2015). For the former, we utilized the xpselect and xpbin applications of ixpeobssim (v31.0.1) to first select photons belonging to the source and background regions and then bin these filtered data files using the PCUBE algorithm (unweighted analysis). This binning routine combines the event-by-event Stokes parameters, subtracts the background contribution, and computes the degree (PD) and angle (PA) of polarization of the source, along with the associated errors, according to the procedure of Kislat et al. (2015). PA represents the electric vector position angle and is conventionally defined anticlockwise relative to the local north (towards northeast). Combining the data from the three IXPE DUs over the 2–8 keV energy range, we obtained with this method a PD for the PWN of oriented at . This PA is similar, although in tension at 3 level, with the measured spin-axis direction of PSR J18331034. We also repeated the same analysis by dividing the data into two energy bins, 2–4 keV and 4–8 keV, and obtained a result that is consistent within uncertainties with a constant polarization throughout the IXPE energy range. The results are summarized on the left side of Table 2 and visualized in Fig. 2.
@@
For the spectro-polarimetric analysis, instead, we started from the same filtered data, produced earlier for the source and background regions, and used the xpbin application (algorithms PHA1, PHA1Q and PHA1U, unweighted analysis) to compute the Stokes I, Q, and U spectra for the three DUs. These files can then be jointly fit in 3ML or xspec444https://heasarc.gsfc.nasa.gov/xanadu/xspec/ to simultaneously recover the spectral and polarimetric parameters of the user-defined models. For G21.50.9, we modeled the observed spectrum with an absorbed power-law, with absorption given by the TbAbs555https://astromodels.readthedocs.io/en/latest/notebooks/TbAbs.html model with wilm abundances (Wilms et al., 2000). According to what was found by Guest et al. (2019), the column density parameter was fixed to the value . Regarding the polarimetric model, based on the result of the previous PCUBE analysis, we chose to employ a model with a constant degree and angle of polarization. To account for potential cross-calibration uncertainties among the different IXPE telescopes, we also added and left free to vary a multiplicative normalization constant for DU2 and DU3 relative to DU1. The best-fit results obtained using 3ML are provided on the right side of Table 2. The power-law index we find is , which is somewhat larger than expectations based on the work by Guest et al. (2019), most likely due to contamination from the much softer external halo and to a different size of the source region. Please note that the uncertainty quoted here represents the statistical error only and does not include any systematics due to potential calibration issues (see Appendix A for further discussion). The polarimetric results agree well with the PCUBE analysis with a best-fit PD = and PA = . Fig. 2 gives a summary of the main results of the space-integrated polarimetric analysis of the PWN.
@@ 3.2 Space-resolved analysis
It has been well known since the beginning of the IXPE mission that any imperfections in the reconstruction of the photon absorption points inside the GPD, due to the event reconstruction algorithm (Baldini et al., 2021) or the spatial resolution of the GPD itself, can cause the appearance of a radially polarized halo around the source, commonly referred to as polarization leakage (for a detailed overview of the effect, see Bucciantini et al. (2023a)). Being an intrinsic effect of the reconstruction algorithm or the GPD, this phenomenon is present in all IXPE observations, but in practice it is only really relevant for the analysis of a handful of bright extended sources, which exhibit a very sharp intensity gradient. Since the polarization pattern induced by the leakage is radial, its contribution effectively averages out when selecting a circular region that includes the entire source, as done in Sect. 3.1. For the space-resolved polarization analysis, instead, polarization leakage can strongly affect the results, and a dedicated analysis is required to estimate and subtract its contribution.
@@
Given the morphology of G21.50.9, and its small relative size comparable to the IXPE angular resolution, it is expected that the effect of polarization leakage will be particularly relevant in the outskirts of the PWN. Its contribution was evaluated using different techniques, including a full ixpesim/ixpeobssim (Di Lalla, 2019) simulation of the source using GEANT4 (Agostinelli et al., 2003), the Mueller matrix formalism described in Bucciantini et al. (2023a), a hybrid machine learning and analytic track reconstruction method (Cibrario et al., 2023) and, finally, the recently released LeakageLib package (Dinsmore & Romani, 2024). The latter tool, in particular, implements the most accurate model of the polarization leakage to date using in-flight calibrated 2D PSFs, and its routines allow the user to predict the leakage pattern based on a Chandra image and subtract it from the measured IXPE polarization map. A detailed summary of the different approaches can be found in Appendix B.
@@
Figure 3 shows the detected IXPE polarization pattern, overlaid on top of the deep Chandra count map of G21.50.9 zoomed over the PWN region. The effect of polarization leakage has been estimated and subtracted with LeakageLib, using the same Chandra image to model the source. The measured polarization information is encoded in a set of segments, with a length proportional to the measured PD, oriented according to the PA (measured east of north) and with a color that depends on the level of significance of the detection: green lines represent measurements with significance , while the black color corresponds to significance between and . The white lines represent the polarization pattern measured in radio (Lai et al., 2022). Measurements consistent with no polarization within are not shown. As can be seen in Fig. 3, the detected pattern in the central region of the PWN, where the detection is most statistically significant, is roughly consistent with a constant PD of 11–12% at an angle of 30–35°. The detection in the outer regions has generally lower significance and the measured PD and PA values vary slightly across the nebula. This fact explains why the space-integrated analysis (Sect. 3.1), which encompasses the whole PWN, measured a lower degree of polarization. The only notable exception is the NW region (top right) of the PWN, where the degree of polarization becomes progressively higher up to a PD of 20–30% and an orientation around 20–25°. No significant polarization was detected outside the PWN.

### v2

@@ 1 Introduction
G21.50.9 is a young, composite plerionic supernova remnant (SNR) characterized in X-rays by the presence of a central, bright, 40″ radius PWN surrounded by a much dimmer and softer SNR shell (Slane et al., 2000; Safi-Harb et al., 2001). Deep observations obtained with Chandra X-ray Observatory (CXO) and XMM-Newton have revealed that the SNR blast wave extends up to 150″ and displays a limb brightening feature at the eastern boundary and knots of enhanced soft X-ray emission above the PWN in the northern direction (Matheson & Safi-Harb, 2005; Bocchino et al., 2005; Matheson & Safi-Harb, 2010; Guest et al., 2019). Observations of G21.50.9 across various electromagnetic wavelengths, from radio to -rays, have unveiled a wealth of information about this source. Using the Very Large Array (VLA) to measure its expansion speed, Bietenholz & Bartel (2008) estimated its age to be yr, making G21.50.9 one of the youngest PWNe known in our galaxy. Its distance to Earth was measured by several authors using different techniques and was found to be around 5 kpc (Safi-Harb et al., 2001; Camilo et al., 2006; Bietenholz & Bartel, 2008; Tian & Leahy, 2008). The first maps at radio wavelengths go back to the 1970–80s (Wilson & Weiler, 1976; Becker & Kundu, 1976; Furst et al., 1988), but even more recent deep searches have failed to detect a radio counterpart of the X-ray SNR shell, with the exception of the northern knot (Bietenholz et al., 2011).
At the heart of G21.50.9 lies the pulsar PSR J18331034, first independently discovered in radio by Gupta et al. (2005) and Camilo et al. (2006), who found a compact pulsating source at the center of the PWN (R.A. = , Dec. = (J2000.0)) with period ms, , surface magnetic field of G, characteristic age of 4.8 kyr, and a spin-down luminosity of . The pulsar remains currently undetected at X-ray energies, while -ray pulsation has been detected with the Fermi Large Area Telescope (Abdo et al., 2013; Smith et al., 2023). Past X-ray observations have only led to upper limits on the pulsating fraction, while still detecting a central compact source surrounded by an elliptical emission region (Slane et al., 2000; Safi-Harb et al., 2001; Camilo et al., 2006). The projected spin-axis direction of PSR J18331034, inferred by fitting the pulsar wind torus using CXO data, was determined to be at a position angle of north through east (Ng & Romani, 2008).
@@
The initial radio polarization maps of G21.50.9 were obtained by Becker & Szymkowiak (1981) with the VLA and by Furst et al. (1988) using the NRO Millimeter-wave Array and revealed a highly ordered, linearly polarized emission distributed in a ring structure, indicating a radial magnetic field configuration. The level of polarization was found to have a minimum towards the center, where little polarization was detected, and to increase with radius up to 20–30% near the boundary of the PWN. Infrared polarimetric observations performed with the Very Large Telescope on the compact core of the PWN (angular size 5″) revealed a highly polarized emission, with an average value of the linear polarization degree of , and a swing of the polarization angle across the inner nebula consistent with a toroidal magnetic field (Zajczyk et al., 2012). On the other hand, a recent polarimetric study by Lai et al. (2022), using archival high-resolution radio observations taken with VLA, confirmed the global radial magnetic field structure of the PWN and ruled out the existence of any strong large-scale toroidal field components extending beyond the inner nebula.
@@ 2 Observations
IXPE observed G21.50.9 in three separate segments close in time and combined in a single observation with a total livetime of ks: the first segment lasted from October 5, 2023 at 00:05:39.233 UTC to October 9, 2023 at 02:16:33.292 UTC, the second one from October 10, 2023 at 22:45:54.503 UTC to October 17, 2023 at 16:42:55.034 UTC and the last one from October 19, 2023 at 21:35:44.584 UTC to October 27, 2023 at 15:39:00.534 UTC. IXPE Level-2 data, publicly available for download on the HEASARC archive111https://heasarc.gsfc.nasa.gov/docs/ixpe/archive/ (ObsID 02001199), were reprocessed and analyzed according to the standard technique using the version 20240125 of the IXPE Instrument Response Functions (IRFs) released on the CALDB222https://heasarc.gsfc.nasa.gov/docs/ixpe/caldb/ on February 28, 2024 (validity date 20230702). Given the temporal proximity and relatively short total duration of the IXPE observations, spanning less than a month, and the known variability timescale of the nebula, which is on the order of several months to years (Guest et al., 2019), G21.50.9 can be considered effectively steady for the purpose of this analysis.
Before analyzing the data, we performed an initial background rejection following the procedure described in Di Marco et al. (2023) to reduce the instrumental background due to cosmic-ray particles interacting in the GPD by approximately 40%. We also excluded approximately 4 ks of total livetime () affected by sudden spikes in the counting rate over short time intervals, likely associated with solar activity and the boundary of the South Atlantic Anomaly. In addition, we corrected the absolute sky coordinates by adjusting the WCS keywords in the FITS header to align the reference frame with that of CXO (TCRPX7 = 304, TCRPX8 = 301.5). These are the data files used from now on in this paper. Fig. 1 shows the count map resulting from the combination of the three IXPE telescopes over the full energy range (2–8 keV). The white circular region (radius 0.8′ centered on the pulsar location) is used to select the inner PWN, while the yellow annulus (inner/outer radii of 3′ and 4′ from the pulsar) defines the background extraction region used for the space-integrated analysis described in Sect. 3.1.
G21.50.9 is observed with CXO routinely as a calibration source. Overlaid in green in Fig. 1 are the contours of a deep image created by merging 18 individual observations listed in Table 1 (total exposure ks) and contained in the Chandra Data Collection https://doi.org/10.25574/cdc.425 (catalog DOI:10.25574/cdc.425), all of which had the remnant placed in the ACIS-S detector, at the default aimpoint of the S3 CCD chip. The events were combined with the merge_obs script in the ciao333https://cxc.cfa.harvard.edu/ciao/ analysis software. A count map was then created using events in the 0.5–8.0 keV energy band, and adaptively smoothed with the dmimgadapt tool using a cone filter with minimum/maximum smoothing radii of 0.5″ and 150″ spanning 300 logarithmically spaced scales, with a minimum of 75 counts within the smoothing kernel. The resulting map is used as background image in all polarization maps shown throughout the paper.
@@
| 159 | 14.85 | 3700 | 9.54 |
| 1230 | 14.56 | 5159 | 9.83 |
| 1433 | 14.97 | 5166 | 10.02 |
| 1554 | 9.06 | 6071 | 9.64 |
| 1717 | 7.54 | 6741 | 9.83 |
| 1770 | 7.22 | 8372 | 10.01 |
| 1838 | 7.85 | 10646 | 9.54 |
| 2873 | 9.83 | 14263 | 9.57 |
| 3693 | 9.78 | 16420 | 9.57 |
Note. — List of ObsIDs and their exposure for the CXO observations used to produce the count map used as background in all polarization maps shown in the paper (e.g. Fig. 3) and whose contours are shown in Fig. 1. The total exposure is around 183 ks. The CXO datasets are contained in the Chandra Data Collection https://doi.org/10.25574/cdc.425 (catalog DOI:10.25574/cdc.425).
@@ 3.1 Space-integrated analysis
As is typically done for other IXPE observations (Bucciantini et al., 2023b; Xie et al., 2022; Negro et al., 2023), using the spatial regions described at the end of Sect. 2 and shown in Fig. 1, we studied the polarization of the entire PWN in two different ways: a model-independent polarimetric analysis using ixpeobssim (Baldini et al., 2022) and a spectro-polarimetric fit with the Multi-Mission Maximum Likelihood (3ML) framework (Vianello et al., 2015). For the former, we utilized the xpselect and xpbin applications of ixpeobssim (v31.0.1) to first select photons belonging to the source and background regions and then bin these filtered data files using the PCUBE algorithm (unweighted analysis). This binning routine combines the event-by-event Stokes parameters, subtracts the background contribution, and computes the degree (PD) and angle (PA) of polarization of the source, along with the associated errors, according to the procedure of Kislat et al. (2015). PA represents the electric vector position angle and is conventionally defined anticlockwise relative to the local north (towards northeast). Combining the data from the three IXPE DUs over the 2–8 keV energy range, we measured with this method a PD of for the PWN, with a PA of . The corresponding Minimum Detectable Polarization at 99% confidence is (Weisskopf et al., 2010), well below the detected level polarization. The measured PA is broadly consistent with the spin-axis direction of PSR J18331034, although in tension at 3 level. To investigate possible energy dependence, we repeated the analysis in two energy bands, 2–4 keV and 4–8 keV, and found no significant variation in polarization, with results consistent within uncertainties. As an internal consistency check, we also repeated the analysis independently for each DU and found agreement within for both PD and PA. The results for the combined DUs are summarized on the left side of Table 2 and visualized in Fig. 2.
@@
For the spectro-polarimetric analysis, instead, we started from the same filtered data, produced earlier for the source and background regions, and used the xpbin application (algorithms PHA1, PHA1Q and PHA1U, unweighted analysis) to compute the Stokes I, Q, and U spectra for the three DUs. These files can then be jointly fit in 3ML or xspec444https://heasarc.gsfc.nasa.gov/xanadu/xspec/ to simultaneously recover the spectral and polarimetric parameters of the user-defined models. For G21.50.9, we modeled the observed spectrum with an absorbed power-law, with absorption given by the TbAbs555https://astromodels.readthedocs.io/en/latest/notebooks/TbAbs.html model with wilm abundances (Wilms et al., 2000). According to what was found by Guest et al. (2019), the column density parameter was fixed to the value . Regarding the polarimetric model, based on the result of the previous PCUBE analysis, we chose to employ a model with a constant degree and angle of polarization. To account for potential cross-calibration uncertainties among the different IXPE telescopes, we also added and left free to vary a multiplicative normalization constant for DU2 and DU3 relative to DU1. The best-fit results obtained using 3ML are provided on the right side of Table 2. The power-law index we measured is , slightly higher than the value reported by Guest et al. (2019), most likely due to contamination from the much softer external halo and differences in the source extraction region. In support of this interpretation, our spectral result is in good agreement with the analysis by Tsujimoto et al. (2011), which used a much larger extraction radius (165″) encompassing the entire remnant. Note that the quoted uncertainty represents the statistical error only and does not account for possible systematic effects from instrumental calibration (see Appendix A for further discussion). The unabsorbed flux (2–8 keV) is . The polarimetric results agree well with those from the PCUBE analysis, with a best-fit PD = and PA = . Fig. 2 gives a summary of the main results of the space-integrated polarimetric analysis of the PWN.
@@ 3.2 Space-resolved analysis
Since the start of the IXPE mission, it has been well known that imperfections in reconstructing photon absorption points within the GPD, arising from the event reconstruction algorithm (Baldini et al., 2021) or from the finite spatial resolution of the GPD, can produce a radially polarized halo around the source, commonly referred to as polarization leakage (for a detailed overview of the effect, see Bucciantini et al. (2023a)). Being an intrinsic effect of the detector or its reconstruction algorithm, this phenomenon is present in all IXPE observations, but in practice it is only really relevant for the analysis of a handful of bright extended sources, which exhibit a very sharp intensity gradient. Since the polarization pattern induced by the leakage is radial, its contribution effectively averages out when selecting a circular region that includes the entire source, as done in Sect. 3.1. For the space-resolved polarization analysis, instead, polarization leakage can strongly affect the results, and a dedicated analysis is required to estimate and subtract its contribution.
@@
Given the morphology of G21.50.9, and its small relative size comparable to the IXPE angular resolution, it is expected that the effect of polarization leakage will be particularly relevant in the outskirts of the PWN. Its contribution was evaluated using different techniques, including a full ixpesim/ixpeobssim (Di Lalla, 2019) simulation of the source using GEANT4 (Agostinelli et al., 2003), the Mueller matrix formalism described in Bucciantini et al. (2023a), a hybrid machine learning and analytic track reconstruction method (Cibrario et al., 2023) and, finally, the recently released LeakageLib package (Dinsmore & Romani, 2024). The latter tool, in particular, implements the most accurate model of the polarization leakage to date using in-flight calibrated 2D PSFs, and its routines allow the user to predict the leakage pattern based on a CXO image and subtract it from the measured IXPE polarization map. A detailed summary of the different approaches can be found in Appendix B. Despite methodological differences, all four techniques yield consistent polarization structures within uncertainties, confirming the robustness of the analysis, particularly in the central, high-significance region of the PWN.
@@
Figure 3 shows the detected IXPE polarization pattern, overlaid on top of the deep CXO count map of G21.50.9 zoomed over the PWN region. The effect of polarization leakage has been estimated and subtracted with LeakageLib, using the same CXO image to model the source. The measured polarization is visualized using a set of segments, one per pixel, whose length and inclination represent the local PD (in scale) and PA (measured east of north). The segment color indicates the level of significance of the detection: green corresponds to measurements with significance , black to those between and . For comparison, white segments represent the polarization pattern measured in the radio band (Lai et al., 2022). Measurements with polarization consistent with zero within are not shown. As shown in Fig. 3, the detected pattern in the central region of the PWN, where the statistical significance is highest, is roughly consistent with a constant PD of 11–12% at an angle of 30–35°. In the outer regions, the detection is generally less significant due to lower photon statistics. While the measured PD and PA values in these regions appear to vary slightly across the nebula, they are most likely still consistent with a uniform polarization pattern when accounting for their larger uncertainties and the systematics introduced by the leakage correction. This apparent variation, along with the inclusion of regions where no significant polarization was observed—such as the SW portion of the PWN (bottom right)— helps explain why the space-integrated analysis (Sect. 3.1), which includes the entire PWN, yielded a lower average PD. The only notable exception is the NW region (top right), where the PD increases progressively, reaching values as high as 20–30% with a PA around 20–25°. No significant polarization was detected outside the PWN.